OpenStudy (anonymous):

trying to find x and y intercepts of 9x^2+4y^2=36 .... i did 9x^2+4(0)^2=36 and i got 9x^2=36 but i dont know what to do from there!

OpenStudy (anonymous):

You pretty much isolate the x and y. So 9x^2 =36 x^2 =4 x=sqrt(4)=2 So the x intercept is ±2

OpenStudy (anonymous):

or (±2,0)

OpenStudy (anonymous):

so you just square the x^2 and the 4?